OpenStudy (anonymous):

Evaluate the expression below when x = 9. 3x + 2x2 - 9 A) 54 B) 180 C) 342 D) 4,365

OpenStudy (anonymous):

Never mind I got it, it is 180 thank you though! @tbl

OpenStudy (anonymous):

you must replace the x with 9 so it would become 3(9) + 2(9)^2 - 9 then you'd mutiply the parathethesis, which would make it 27 + 18^2 - 9 then you would do the exponets which would mean 18 * 18 which would make it 324 the problem would then change to 27 + 324 - 9 you would then do the logical thing add and subtract. 27 + 324 which would become 351 then you'd minus 9 which would make it 342 which would make your answer C

OpenStudy (anonymous):

oh I got 180

OpenStudy (anonymous):

erm how?

OpenStudy (anonymous):

would you like explain your steps?

OpenStudy (anonymous):

herm i don't see how you could get 180...

OpenStudy (anonymous):

is that x squared, or times 2?

OpenStudy (anonymous):

haha I don't know I think I must have added or subtracted it wrong but already put it in and cant go back thank you though @tbl

OpenStudy (anonymous):

its x squared

OpenStudy (anonymous):

rawr should have let me finished tina! but good try!

OpenStudy (anonymous):

\[3x + 2x ^{2} - 9 = Z\] \[3(9) + 2(9) ^{2} - 9 = Z\] \[27+ 324 - 9 = Z\] \[351 - 9 = Z\] \[342 = Z\]

OpenStudy (anonymous):

what's with the z and that's what i said already

OpenStudy (anonymous):

thank you! @RockerSk8er I see my mistake.

OpenStudy (anonymous):

guess my explnation didn't help but good job @RockerSk8er

OpenStudy (anonymous):

lol I meant it as you already explained it to me and I understood it already! @tbl

OpenStudy (anonymous):

thank you both though ♥ @tbl @RockerSk8er

OpenStudy (anonymous):

welcome tina! and any time you need help just tag me in

OpenStudy (anonymous):

I just added Z to represent the final number. Not that yours didn't say it all, I just put it into an expression. :) You're welcome @tinasanchez

OpenStudy (anonymous):

ok thanks @tbl

OpenStudy (anonymous):

erm your welcome..:)?